Melinda has nickels and quarters in her bank. She has seven fewer nickels than quarters. She has $2.65 in the bank. How many coins of each type does she have?
Answer by addingup(3677)

q = n+7
0.05n+0.25q = 2.65 substitute for q:
0.05n+0.25(n+7) = 2.65
0.05n+0.25n+1.75 = 2.65
0.30n = 0.90
n = 3 she has 3 nickels and
2.65-0.15= 2.50 = 10 quarters